Two items will be selected at random from a pencil box. The pencil box contains 18 items: 4pencils, 10 pens, and 4 markers. The first item will be selected at random; then the second item will be selected randomly from the remaining items in the pencil box. What is the probability that two pens will be selected? Report your answer as an exact fraction. Simplify your answer as much as possible.
Answer by venugopalramana(3286)

Two items will be selected at random from a pencil box. The pencil box contains 18 items: 4pencils, 10 pens, and 4 markers. The first item will be selected at random; then the second item will be selected randomly from the remaining items in the pencil box. What is the probability that two pens will be selected? Report your answer as an exact fraction. Simplify your answer as much as possible.
THERE ARE 10 PENS OUT OF TOTAL 18 ITEMS
HENCE PROBABILITY OF PICKING A PEN ON I ATTEMPT=10/18=5/9
NOW HAVING ACHIEVED ABOVE ,THERE ARE 9 PENS REMAINING OUT OF 17 ITEMS
HENCE PROBABILITY OF PICKING A PEN ON II ATTEMPT=9/17
SO COMBINED PROBABILITY OF PICKING 2 PENS IS =(5/9)*(9/17)=5/17=0.294